Lenght 28 cm and width 17 cm. What is the area and perimeter?
Answer by aditya2117(32)

Area = 28 * 17 sq.cm
= 476 sq.cm
Perimeter = 2(28+17) cm
= 2*45 cm
= 90cm
That's It!!